OpenStudy (anonymous):

Integrals PLEASE HELP what is the integral of e/x^(4)+e^(2)/sqrt(x)dx

OpenStudy (anonymous):

So hopefully I read that right: \[\int\limits_{}^{}(\frac{e}{x^4}+\frac{e^2}{\sqrt{x}})dx\]This can be split up because the integral of a sum is the sum of the integrals. \[\int\limits \frac{e}{x^4}dx+\int\limits \frac{e^{2}}{\sqrt{x}}dx\] Dividing by x^4 is just x^-4, square root is just 1/2 power, and the e and e^2 are just constants so they may be pulled out. \[e \int\limits x^{-4}dx + e^{2}\int\limits x^{1/2}dx \] Now it's just simple power rule and this is the answer: \[e \frac{x^{-3}}{-3} + e^{2}\frac{x^{3/2}}{3/2}=e(\frac{2}{3}ex^{3/2}-\frac{1}{3x^{3}})\]

OpenStudy (anonymous):

I'm sorry, that's supposed to be an x^(-1/2), it's not supposed to be postive. Which means all the 3/2 are supposed to be 1/2 and the 2/3 should be a 2. Is there any way to edit my answer?